# Non-transitive pseudo-Anosov flows

### v1

@@ 4.3. Smale Chains
Suppose $x,y\in\mathcal{R}_{G}$, $x\nsim_{G}y$, and $\mathcal{F}^{+}(x)\cap\mathcal{F}^{-}(y)\neq\emptyset$. Then there exists a Smale chain $\mathcal{W}$ such that $\mathcal{F}^{+}(x)\cap\mathcal{F}^{-}(y)$ is contained in $\mathcal{W}$ and the closure of the Smale class of $x$ contains a corner of $\mathcal{W}$.
@@ 5. Virtually cyclic point stabilizers and $\mathbb{Z}^{2}$ invariant Smale chains
Theorem 1.4 does not hold in the $\mathbb{R}$-covered setting. Example 1 of [BFM22] gives Anosov-like affine actions on trivial planes with indiscrete stabilizers. On skew planes, one may create such examples by taking subgroups of $\mathrm{PSL}(2,\mathbb{R})$ with non-discrete point stabilizers (but no parabolic elements), and lifting them to $\widetilde{\mathrm{PSL}}(2,\mathbb{R})$. The action of $\widetilde{\mathrm{PSL}}(2,\mathbb{R})$ induces an action on a skew plane, and such a subgroup without parabolics will induce an Anosov-like action on the plane.
@@
Let $H_{x}$ denote the finite index subgroup of the stabilizer of $x$ fixing each ray of ±$\mathcal{F}^{\pm}(x)$. If $H_{x}$ is not cyclic, then for any point $p$ in any TL subquadrant of $x$, the orbit $H_{x}(p)$ has an accumulation point.
@@
Suppose first that $G$ acts topologically transitively on $P$. Since $P$ is not $\mathbb{R}$-covered, there is a branching leaf or prong. By [BFM22, Lemma 2.21], every leaf of $\mathcal{F}^{\pm}$ has a dense orbit under $G$. Thus, some branching leaf or prong leaf must intersect $r^{+}$, and Claim 5.3 applies to give the conclusion.
Now suppose instead that $G$ has a non-transitive Anosov like action. We first make a short argument that will allow us to treat only the case where $x$ is nonsingular. Recall that $x$ is fixed by some element, and $Q$, the $TL$-subquadrant bounded by $r^{\pm}$ was assumed not to be a lozenge. Hence it is a regular subquadrant, and therefore we can find $x^{\prime}$ a nonsingular point in $Q$ totally linked with $x$ and such that $x^{\prime}$ is fixed by some nontrivial element $g^{\prime}$ of $G$. Hence if we can find $y\in\mathrm{Pivot}\cup\mathrm{Sing}$ such that $\mathcal{F}^{+}(y)\cap\mathcal{F}^{-}(x^{\prime})\neq\emptyset$ (resp. $\mathcal{F}^{-}(y)\cap\mathcal{F}^{+}(x^{\prime})\neq\emptyset$) then for some large enough power $n$, we have $\mathcal{F}^{+}((g^{\prime})^{n}y)\cap r^{-}\neq\emptyset$ (resp. $\mathcal{F}^{-}((g^{\prime})^{n}y)\cap r^{+}\neq\emptyset$).
Thus we can now assume that $x$ is nonsingular and just find an element of $y\in\mathrm{Pivot}\cup\mathrm{Sing}$ such that at least one of ±$\mathcal{F}^{\pm}(y)$ intersects $\mathcal{F}^{\mp}(x)$.
Since we assumed the action of $G$ was non-transitive, by Theorem 1.8, there exists an element $y^{\prime}$ in a distinct Smale class from $x$. By Theorem 4.23, there exists a Smale chain $\mathcal{W}$ intersecting either $\mathcal{F}^{+}(x)$ of $\mathcal{F}^{-}(x)$ (or, in general, if $x$ is neither a maximal nor minimal Smale class, then we will find examples of both). In particular, there exists a corner $y$ on $\mathcal{W}$ with ±$\mathcal{F}^{\pm}(y)\cap\mathcal{F}^{\mp}(x)\neq\emptyset$. Since all corners of a Smale chain are either pivots or singular points, the hypothesis for Claim 5.3 holds, and implies our claim. ∎
@@
We now assume that we are in case 2: $L_{n}$ intersects the ray $r^{-}$, but not $r^{+}$. That is, $L_{n}$ is in the top left quadrant of $y_{n}$. As above, up to renaming the $y_{n}$, we may assume that $y_{1}=gy_{0}$ is to the bottom right of $y_{0}$ and totally linked with it. But this forces $z_{1}$, which is a pivot or a prong to be inside $L_{0}$, which is impossible (by the “non-corner criterion” [BFM22, Lemma 2.29], as in the argument of the proof of Lemma 2.21).
@@ 5.1. $\mathbb{Z}^{2}$ invariant Smale chains
Let $\mathcal{W}:=\{L_{i},i\in I\}$ and $\mathcal{W}^{\prime}:=\{L^{\prime}_{i},i\in I^{\prime}\}$ be two chains of lozenges (finite or infinite, with index sets $I$ and $I^{\prime}$ some consecutive subsets of integers containing 00), satisfying conditions (i)(iii) given in Definition 4.20. Assume that there exists $k$ such that, for $0\leq i\leq k$, $L_{i}=L^{\prime}_{i}$. If $k=0$, suppose moreover that the shared corner of $L_{-1}$ and $L_{0}$ is the same as the shared corner of $L^{\prime}_{-1}$ and $L^{\prime}_{0}=L_{0}$. Then $\mathcal{W}^{\prime\prime}:=\{L_{i},i\leq k\}\cup\{L^{\prime}_{i},i\geq k+1\}$ also satisfies conditions (i)(iii).
@@
6. Density of non-corner leaves
@@
A non-corner fixed leaf is a leaf of $\mathcal{F}^{\pm}$ that does not contain a corner of a lozenge but is fixed by some nontrivial element of $g$. Equivalently, it is a leaf of a non-corner fixed point. Axiom (A2) says that the union of leaves (of either foliation) containing fixed points is dense. The point of this section is to show density of non-corner fixed leaves for Smale-bounded actions. This will be proved in two steps, first we will show density on extremal Smale classes, and then use Smale-boundedness to deduce that property on the whole plane.
@@
Before we start with the proof, we will prove one more preliminary result that will help us simplify the argument. This is a refinement of the fact that being a corner point is a closed property among points fixed by nontrivial elements in $G$ [^7] Notice that it may not be a closed property in general, see Proposition 3.5 of [Fen22].
@@
Up to taking a subsequence, we may assume that all $x_{n}$ are pairwise totally linked, totally linked with $x$, and nonsingular. We may further assume that all $x_{n}$ are in the same quadrant $Q$ of $x$ and that their leaves converge monotonically towards those of $x$ and hence all the $x_{n}$ lie in the product foliated rectangle defined by ±$\mathcal{F}^{\pm}(x_{1})$, ±$\mathcal{F}^{\pm}(x)$ and their intersections. Calling $y_{n}$ the other corner of $L_{n}$, we can also assume that the $y_{n}$ are all in the same quadrant $Q^{\prime}$ of $x$. (The quadrant $Q^{\prime}$ could be the same as $Q$, or one of the two adjacent ones, or possibly, if $x$ is nonsingular, the opposite quadrant). This implies that, for any $n$ we have $L_{n}\cap L_{1}\neq\emptyset$.
@@
Figure 12. .
@@
Thus, since there are no infinite product regions, $\mathcal{F}^{+}(y_{n})$ must converge to a single leaf, or to a union of nonseparated leaves. Since $\mathcal{F}^{+}(y_{n})$ makes a perfect fit with $\mathcal{F}^{-}(x_{n})$, we have that one leaf in the limit makes a perfect fit with $\mathcal{F}^{-}(x)$. Since $x$ was assumed to be fixed by $g$, and one of its leaves makes a perfect fit, Lemma 2.17 implies that $x$ is the corner of a lozenge $L$. By construction, one side of $L$ through $x$ is in the limit of the side of $L_{n}$, one easily deduce that the other side of $L$ through $x$ is also in the limit of a side of $L_{n}$. ∎
We also make the following observation, which generalizes the well-know fact that hyperbolic attractors (resp. repellor) of an Anosov flow are saturated by unstable (resp. stable) leaves.
Observation 6.3.
@@
For concreteness, assume that $\Lambda$ is maximal rather than minimal. If $\Lambda$ is the unique Smale class, then, by Theorem 1.8, $G$ is a transitive Anosov-like action in the sense of [BFM22], so the non-corner fixed points are dense by Lemma 2.30 of [BFM22].
@@
By Theorem 4.23, there exists a Smale chain $\mathcal{W}$ with a corner $c\in\bar{\Lambda}$. By Observation 6.3, $\Lambda$ is $\mathcal{F}^{+}$ saturated.
Let $L_{1}$ be a lozenge of $\mathcal{W}$ with corner $c$. Since $\Lambda$ is $\mathcal{F}^{+}$ saturated (Observation 6.3), there exists a face of $\mathcal{F}^{+}(c)$ that is accumulated by points of $\Lambda$; thus $c$ has a pair of regular TL subquadrants, that we denote by $R_{1}$ and $R_{2}$ which are adjacent (and share a ray of $\mathcal{F}^{-}(c)$). Up to switching $L_{1}$ to another wandering subquadrant if needed, we assume $L_{1}$ is adjacent to $R_{1}$. If either $R_{1}$ or $R_{2}$ contains non-corner points of $\mathrm{Fix}_{G}$, these lie in $\Lambda$ and we are done, so we assume such points are all corners.
@@
Suppose $x$ is a corner point in $R_{1}\cap\mathrm{Fix}_{G}$. Lemma 2.29 in [BFM22] says the lozenge cannot contain $c$, due to the perfect fit formed by sides of $L_{1}$. This eliminates one of the two upper quadrants (the top right). To finish the proof, we will show that any small neighborhood of $x$ contains an open set of points of $R_{1}\cap\mathrm{Fix}_{G}$ that are either non-corner, or only have lozenges in their lower quadrant. To do this, consider some nonsingular corner point $y$ in $R_{2}$ totally linked with $x$. By Corollary 3.10, for any trivially foliated neighborhood of $x$, there exists $z\in R_{1}\cap\mathrm{Fix}_{G}$ in that neighborhood and $g\in G$ such that $g(z)$ lies arbitrarily close to $y$, and we can take $g$ to preserve the local orientation. By the first part $z$ cannot have a lozenge in its upper left quadrant. Now, either $c$ is singular, or it is nonsingular and its remaining quadrant is a wandering lozenge. In either case, applying Lemma 2.29 of [BFM22] to $g(z)$ shows that it cannot be the corner of a lozenge $L$ in the quadrant containing $c$. Equivalently, $z$ cannot have a lozenge in its other upper quadrant, proving the claim. ∎
@@
More generally, if for any Smale class $\Lambda_{i}$ there exists $\Lambda_{j}$ maximal with $\Lambda_{i}\leq_{G}\Lambda_{j}$ then the set of non-corner fixed leaves in $\mathcal{F}^{-}$ is dense. Similarly, if for any Smale class $\Lambda_{i}$ there exists $\Lambda_{k}$ minimal with $\Lambda_{k}\leq_{G}\Lambda_{i}$ then the set of non-corner fixed leaves in $\mathcal{F}^{+}$ is dense.
@@
The first claim follows trivially from the second two, which are symmetric up to switching the roles of $\mathcal{F}^{+}$ and $\mathcal{F}^{-}$. So we only prove the density of simple $\mathcal{F}^{+}$ leaves, assuming that all Smale class admits a lower bound.
Let $U$ be some open set in $P$. By Axiom (A2), there exists $x$ fixed by some $g\in G$ with $\mathcal{F}^{+}(x)\cap U\neq\emptyset$. By assumption, there exists a minimal Smale class $\Lambda$ such that $\mathcal{F}^{+}(\Lambda)\cap\mathcal{F}^{-}(x)\neq\emptyset$. Thus, for some high enough power $n$, $g^{n}(U)$ will intersects $\mathcal{F}^{+}(\Lambda)$. By Proposition 6.5 non-corner fixed leaves of $\mathcal{F}^{+}$ are dense in $\mathcal{F}^{+}(\Lambda)$, so there is a non-corner fixed leaf intersecting $g^{n}U$, proving density. ∎

### v2

@@ 4.3. Smale Chains
Suppose $x$ and $y$ are in distinct Smale classes and $\mathcal{F}^{+}(x)\cap\mathcal{F}^{-}(y)\neq\emptyset$. Then there exists a Smale chain $\mathcal{W}$ such that $\mathcal{F}^{+}(x)\cap\mathcal{F}^{-}(y)$ is contained in $\mathcal{W}$ and the closure of the Smale class of $x$ contains a corner of $\mathcal{W}$.
@@
Suppose first that $x$ is an isolated prong singularity. Then by Proposition 4.12, $x$ has a wandering lozenge in each of its quadrants. Thus, $\mathcal{F}^{+}(x)\cap\mathcal{F}^{-}(y)$ is on a side shared by two wandering lozenges with corner $x$. Proposition 4.22 says this pair of lozenges may be extended to a Smale chain, and the conclusion follows.
The same argument applies if $y$ is an isolated prong singularity. So we now assume that both $x$ and $y$ are in $\mathcal{R}_{G}$.
@@ 5. Virtually cyclic point stabilizers and $\mathbb{Z}^{2}$ invariant Smale chains
Theorem 1.2 does not hold in the $\mathbb{R}$-covered setting. Example 1 of [BFM25b] gives Anosov-like affine actions on trivial planes with indiscrete stabilizers. On skew planes, one may create such examples by taking subgroups of $\mathrm{PSL}(2,\mathbb{R})$ with non-discrete point stabilizers (but no parabolic elements), and lifting them to $\widetilde{\mathrm{PSL}}(2,\mathbb{R})$. The action of $\widetilde{\mathrm{PSL}}(2,\mathbb{R})$ induces an action on a skew plane, and such a subgroup without parabolics will induce an Anosov-like action on the plane.
@@
Let $H_{x}$ denote the finite index subgroup of the stabilizer of $x$ fixing each ray of $\mathcal{F}^{\pm}(x)$. If $H_{x}$ is not cyclic, then for any point $p$ in any TL subquadrant of $x$, the orbit $H_{x}(p)$ has an accumulation point.
@@
Suppose first that $G$ acts topologically transitively on $P$. Since $P$ is not $\mathbb{R}$-covered, there is a branching leaf or prong. By [BFM25b, Lemma 2.21], every leaf of $\mathcal{F}^{\pm}$ has a dense orbit under $G$. Thus, some branching leaf or prong leaf must intersect $r^{+}$, and Claim 5.3 applies to give the conclusion.
Now suppose instead that $G$ has a nontransitive Anosov like action. We first make a short argument that will allow us to treat only the case where $x$ is nonsingular. Recall that $x$ is fixed by some element, and $Q$, the $TL$-subquadrant bounded by $r^{\pm}$ was assumed not to be a lozenge. Hence it is a regular subquadrant, and therefore we can find $x^{\prime}$ a nonsingular point in $Q$ totally linked with $x$ and such that $x^{\prime}$ is fixed by some nontrivial element $g^{\prime}$ of $G$. Hence if we can find $y\in\mathrm{Pivot}\cup\mathrm{Sing}$ such that $\mathcal{F}^{+}(y)\cap\mathcal{F}^{-}(x^{\prime})\neq\emptyset$ (resp. $\mathcal{F}^{-}(y)\cap\mathcal{F}^{+}(x^{\prime})\neq\emptyset$) then for some large enough power $n$, we have $\mathcal{F}^{+}((g^{\prime})^{n}y)\cap r^{-}\neq\emptyset$ (resp. $\mathcal{F}^{-}((g^{\prime})^{n}y)\cap r^{+}\neq\emptyset$).
Thus we can now assume that $x$ is nonsingular and just find an element of $y\in\mathrm{Pivot}\cup\mathrm{Sing}$ such that at least one of $\mathcal{F}^{\pm}(y)$ intersects $\mathcal{F}^{\mp}(x)$.
Since we assumed the action of $G$ was nontransitive, by Theorem 1.7, there exists an element $y^{\prime}$ in a distinct Smale class from $x$. By Theorem 4.23, there exists a Smale chain $\mathcal{W}$ intersecting either $\mathcal{F}^{+}(x)$ of $\mathcal{F}^{-}(x)$ (or, in general, if $x$ is neither a maximal nor minimal Smale class, then we will find examples of both). In particular, there exists a corner $y$ on $\mathcal{W}$ with $\mathcal{F}^{\pm}(y)\cap\mathcal{F}^{\mp}(x)\neq\emptyset$. Since all corners of a Smale chain are either pivots or singular points, the hypothesis for Claim 5.3 holds, and implies our claim. ∎
@@
We now assume that we are in case 2: $L_{n}$ intersects the ray $r^{-}$, but not $r^{+}$. That is, $L_{n}$ is in the top left quadrant of $y_{n}$. As above, up to renaming the $y_{n}$, we may assume that $y_{1}=gy_{0}$ is to the bottom right of $y_{0}$ and totally linked with it. But this forces $z_{1}$, which is a pivot or a prong to be inside $L_{0}$, which is impossible (by the “non-corner criterion” [BFM25b, Lemma 2.29], as in the argument of the proof of Lemma 2.23).
@@ 5.1. $\mathbb{Z}^{2}$ invariant Smale chains
Let $\mathcal{W}:=\{L_{i},i\in I\}$ and $\mathcal{W}^{\prime}:=\{L^{\prime}_{i},i\in I^{\prime}\}$ be two chains of lozenges (finite or infinite, with index sets $I$ and $I^{\prime}$ some consecutive subsets of integers containing 0), satisfying conditions (i)(iii) given in Definition 4.20. Assume that there exists $k$ such that, for $0\leq i\leq k$, $L_{i}=L^{\prime}_{i}$. If $k=0$, suppose moreover that the shared corner of $L_{-1}$ and $L_{0}$ is the same as the shared corner of $L^{\prime}_{-1}$ and $L^{\prime}_{0}=L_{0}$. Then $\mathcal{W}^{\prime\prime}:=\{L_{i},i\leq k\}\cup\{L^{\prime}_{i},i\geq k+1\}$ also satisfies conditions (i)(iii).
@@
6. Smale bounded actions and density of non-corner leaves
@@
A non-corner fixed leaf is a leaf of $\mathcal{F}^{\pm}$ that does not contain a corner of a lozenge but is fixed by some nontrivial element of $g$. Equivalently, it is a leaf of a non-corner fixed point. Axiom (A2) says that the union of leaves (of either foliation) containing fixed points is dense. The point of this section is to show density of non-corner fixed leaves under a mild assumption on the actions. This will be proved in two steps, first we will show density on extremal Smale classes, and then use one extra assumption, called Smale-bounded, introduced below, which says that the saturations of extremal Smale classes are dense to deduce the density of non-corner fixed points on the whole plane.
6.1. Smale-bounded actions
Recall (see Definition 3.7) that a Smale class is called extremal if it is regular and either minimal or maximal for the Smale order.
Definition 6.1 (Smale-bounded actions).
We say that an Anosov-like action $G$ is Smale-bounded if the following two conditions hold:
  1. (i)
    For any regular Smale class $\Lambda$, there exist extremal Smale classes $\Lambda_{m},\Lambda_{M}$ such that $\Lambda_{m}\leq_{G}\Lambda\leq_{G}\Lambda_{M}$,
  2. (ii)
    For any singular Smale class $\Lambda$, $p\in\Lambda$, and any face $f^{+}$ of $\mathcal{F}^{+}(p)$, resp. $f^{-}$ of $\mathcal{F}^{-}(p)$, there exists a maximal class $\Lambda_{M}$, resp. minimal class $\Lambda_{m}$, such that $\mathcal{F}^{-}(\Lambda_{M})\cap f^{+}\neq\emptyset$, resp. $\mathcal{F}^{+}(\Lambda_{m})\cap f^{-}\neq\emptyset$.
The point of this definition is just to say that one cannot have an infinite increasing or decreasing sequence of Smale classes, but since the order is not well define on singular Smale class, we have to treat them separately.
We will show that if an action is cocompact, or more generally has only finitely many Smale classes, then it is Smale-bounded. Before doing this, we give the following equivalent definition, which is the property needed to prove density of non-corner fixed points.
Lemma 6.2.
An Anosov-like action $G$ is Smale-bounded if and only if for any open set $U$, there exists a maximal Smale class $\Lambda_{M}$ and a minimal Smale class $\Lambda_{m}$ such that $\mathcal{F}^{-}(\Lambda_{M})\cap U\neq\emptyset$ and $\mathcal{F}^{+}(\Lambda_{m})\cap U\neq\emptyset$.
Proof.
Suppose that $G$ is Smale bounded and consider $U$ an open set. By Axiom (A2), there exists $x$ fixed by some $g\in G$ with $\mathcal{F}^{+}(x)\cap U\neq\emptyset$.
Up to switching $g$ with $g^{-1}$, we have that $g^{n}(U)$ will contain a face $f^{-}$ of $\mathcal{F}^{-}(x)$ in its limit as $n\to+\infty$. By assumption, there exists a minimal Smale class $\Lambda$ such that $\mathcal{F}^{+}(\Lambda)\cap f^{-}\neq\emptyset$. Thus, for some high enough power $n$, $g^{n}(U)$ will intersects $\mathcal{F}^{+}(\Lambda)$ and $G$-invariance of Smale classes yields the result.
The converse direction is obvious: if $x$ is any point in a (regular or singular) Smale class and $f^{\pm}$ is a face of $\mathcal{F}^{\pm}(x)$, then we can take a trivially foliated open set $U$ containing a segment of $f^{\pm}$ in its boundary and the definition of Smale bounded follows immediately. ∎
Next, we want to show that many natural Anosov-like actions (such as those coming from pseudo-Anosov flows on compact 3-manifolds) are Smale-bounded.
Proposition 6.3.
If an Anosov-like action has only finitely many Smale classes, then it is Smale-bounded. In particular, if the action of $G$ on $P$ is co-compact, then it is Smale-bounded.
Proof.
Suppose that $G$ has finitely many Smale classes $\Lambda_{1},\dots,\Lambda_{n}$. By Lemma 3.15, at least one of these classes is regular. By definition, each regular Smale class has minimal Smale class below it and maximal one above it for the Smale order.
Let $U$ be a trivially foliated open set in $P$. By Axiom (A2), there exists $x$ fixed by some $g\in G$ with $\mathcal{F}^{+}(x)\cap U\neq\emptyset$. If $x$ is in a regular Smale class, then a minimal class below the class of $x$ will have a point $p$ with $\mathcal{F}^{+}(p)\cap U\neq\emptyset$. Similarly, we can find a fixed point $y$ with $\mathcal{F}^{-}(y)\cap U\neq\emptyset$ and if $y$ is in a regular class then we have some maximal $\Lambda$ with $\mathcal{F}^{-}(\Lambda)\cap U\neq\emptyset$. Thus, to prove the first statement of the proposition, we need only eliminate the case that the only possibilities for such $x$ (or $y$) are isolated prong singularities. We treat the case of $x$, the other being symmetric.
Suppose for a contradiction that every point $x$ in $\mathrm{Fix}_{G}$ such that $\mathcal{F}^{+}(x)\cap U\neq\emptyset$ is an isolated prong singularity. Let $x_{1}$ be such a point. Then all the quadrants of $x_{1}$ are wandering lozenges (by Proposition 4.12), so there exists a lozenge $L_{1}$ with corner $x_{1}$ such that $L_{1}\cap U\neq\emptyset$. Now pick $x_{2}$ an isolated prong singularity such that $\mathcal{F}^{+}(x_{2})$ intersects $L_{1}\cap U$ and consider, similarly, a lozenge $L_{2}$ that intersects $L_{1}\cap U$. We can keep iterating this process and get a sequence $x_{i}$ of isolated prong singularities and lozenges $L_{i}$ which all intersect. Because there are only finitely many Smale classes, we can extract a subsequence such that all the $x_{i}$ are in the same $G$-orbit. In particular, (as in the proof of Lemma 3.15) we can find $g\in G$ such that $x_{k}=gx_{i}$ for some $i$ and $k$ and such that $L_{k}=gL_{i}$ getting a (nonsingular) fixed point inside $L_{i}$ which is a contradiction.
Finally, we will show that if $G$ acts co-compactly, then the number of Smale classes must be finite: Let $\Lambda_{i}$ be a Smale class. If $\Lambda_{i}$ is regular, then each $x\in\Lambda_{i}$ has a neighborhood $U_{x}$ that is product foliated, so any $y\in U_{x}$ satisfies $y\sim_{G}x$. Suppose $G$ acts cocompactly on $P$. Since Smale classes are $G$-invariant, each is represented in a compact fundamental domain. By the previous remark, there can be no accumulation points of distinct regular Smale classes, and this is also the case by definition for singular Smale classes, so there are only finitely many Smale classes. ∎
We note that there exist example of Anosov-like actions that are not cocompact but are Smale bounded, see Example 8.2. It is also easy to build examples of actions that are not bounded, even in the flows setting, if one allows for non-compactness, see Example 8.4.
6.2. Density of non-corner fixed points
@@
Before we start with the proof, we will prove one more preliminary result that will help us simplify the argument. This is a refinement of the fact that being a corner point is a closed property among points fixed by nontrivial elements in GG. [^9] Notice that it may not be a closed property in general, see Proposition 3.5 of [Fen22].
@@
Up to taking a subsequence, we may assume that all $x_{n}$ are pairwise totally linked, totally linked with $x$, and nonsingular. We may further assume that all $x_{n}$ are in the same quadrant $Q$ of $x$ and that their leaves converge monotonically towards those of $x$ and hence all the $x_{n}$ lie in the product foliated rectangle defined by $\mathcal{F}^{\pm}(x_{1})$, $\mathcal{F}^{\pm}(x)$ and their intersections. Calling $y_{n}$ the other corner of $L_{n}$, we can also assume that the $y_{n}$ are all in the same quadrant $Q^{\prime}$ of $x$. (The quadrant $Q^{\prime}$ could be the same as $Q$, or one of the two adjacent ones, or possibly, if $x$ is nonsingular, the opposite quadrant). This implies that, for any $n$ we have $L_{n}\cap L_{1}\neq\emptyset$.
@@
Figure 12. Possible configurations of sequences of corners
@@
Since there are no infinite product regions, $\mathcal{F}^{+}(y_{n})$ must converge to a single leaf, or to a union of nonseparated leaves. Since $\mathcal{F}^{+}(y_{n})$ makes a perfect fit with $\mathcal{F}^{-}(x_{n})$, we have that one leaf in the limit makes a perfect fit with $\mathcal{F}^{-}(x)$. Since $x$ was assumed to be fixed by $g$, and one of its leaves makes a perfect fit, Lemma 2.19 implies that $x$ is the corner of a lozenge $L$. By construction, one side of $L$ through $x$ is in the limit of the side of $L_{n}$, one easily deduce that the other side of $L$ through $x$ is also in the limit of a side of $L_{n}$. ∎
We also make the following observation, which generalizes the well-know fact that hyperbolic attractors (resp. repellers) of an Anosov flow are saturated by unstable (resp. stable) leaves.
Observation 6.6.
@@
For concreteness, assume that $\Lambda$ is maximal rather than minimal. If $\Lambda$ is the unique Smale class, then, by Theorem 1.7, $G$ is a transitive Anosov-like action in the sense of [BFM25b], so the non-corner fixed points are dense by Lemma 2.30 of [BFM25b].
@@
By Theorem 4.23, there exists a Smale chain $\mathcal{W}$ with a corner $c\in\bar{\Lambda}$. By Observation 6.6, $\Lambda$ is $\mathcal{F}^{+}$ saturated.
Let $L_{1}$ be a lozenge of $\mathcal{W}$ with corner $c$. Since $\Lambda$ is $\mathcal{F}^{+}$ saturated (Observation 6.6), there exists a face of $\mathcal{F}^{+}(c)$ that is accumulated by points of $\Lambda$; thus $c$ has a pair of regular TL subquadrants, that we denote by $R_{1}$ and $R_{2}$ which are adjacent (and share a ray of $\mathcal{F}^{-}(c)$). Up to switching $L_{1}$ to another wandering subquadrant if needed, we assume $L_{1}$ is adjacent to $R_{1}$. If either $R_{1}$ or $R_{2}$ contains non-corner points of $\mathrm{Fix}_{G}$, these lie in $\Lambda$ and we are done, so we assume such points are all corners.
@@
Suppose $x$ is a corner point in $R_{1}\cap\mathrm{Fix}_{G}$. Lemma 2.29 in [BFM25b] says the lozenge cannot contain $c$, due to the perfect fit formed by sides of $L_{1}$. This eliminates one of the two upper quadrants (the top right). To finish the proof, we will show that any small neighborhood of $x$ contains an open set of points of $R_{1}\cap\mathrm{Fix}_{G}$ that are either non-corner, or only have lozenges in their lower quadrant. To do this, consider some nonsingular corner point $y$ in $R_{2}$ totally linked with $x$. By Corollary 3.11, for any trivially foliated neighborhood of $x$, there exists $z\in R_{1}\cap\mathrm{Fix}_{G}$ in that neighborhood and $g\in G$ such that $g(z)$ lies arbitrarily close to $y$, and we can take $g$ to preserve the local orientation. By the first part $z$ cannot have a lozenge in its upper left quadrant. Now, either $c$ is singular, or it is nonsingular and its remaining quadrant is a wandering lozenge. In either case, applying Lemma 2.29 of [BFM25b] to $g(z)$ shows that it cannot be the corner of a lozenge $L$ in the quadrant containing $c$. Equivalently, $z$ cannot have a lozenge in its other upper quadrant, proving the claim. ∎
@@
By Lemma 6.2, the $\mathcal{F}^{+}$-saturation of the union of minimal Smale classes is dense in $P$, and so is the $\mathcal{F}^{-}$-saturation of the union of maximal Smale classes. The result follows then trivially from Proposition 6.8. ∎